INAF queue observing, 2018 November 09-10

Observer: R. Carini, M. Faccini
Telescope Operator: D. Gonzal-Huerta
Support Astronomer: O. Kuhn
Instrument: MODS1, LUCI2

Summary

Night duration (12\xB0 el twilight) 1:15-12:45 UT = 11h 30m

Observing time: 8h 0min (70%)

Weather time loss: 3h 0m (26%)

technical time loss 0h 30m (4%)

Data Summary

At the sunset the sky is mostly clear, wind between 5 - 10 m/s (front - rear) , humidity about 60%

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
Testa A   MODS 1 ls dual grating 0.97/3 h mods1r.0005-0007,mods1b.0001-0003 the second and third images have seeing out of specific
Antoniucci V733 Cep mods1 ls dual grating 0.5/0.5 mods1r.0011-0014, mods1b.0004-0007 completed
standard BD+28 mods1 ls red and dual grating   mods1r.0020-0022,mods1b.0008-0010  
Fabrizio shall1 Tri II mods1 mos dual grating 0.5/0.5 mods1r.0026,mods1b.0011 completed
Fabrizio shall2 Tri II mods1 mos dual grating 0.5/0.5 mods1r.0030,mods1b.0012 completed
Perna telluric HIP16111 luci2 mos   zjspec luci2.0029-0032  
Perna RCSGA327 luci2 mos zjspec 0.6/3.0 h luci2.0039-0053 completed
Nisini field 3 luci2 ima J, H, Ks 0.1 luci2.0054-0079 completed
Pinna field3 HD16760 luci2 AO Ks luci2.0080-0088  

Details

Calibration MODS
  • bias mods1r.20181109.0001-0030,mods1b.20181109.0001-0030

Calibration LUCI 1 & 2
  • dark
  • 1X160 luci1.20181110.0001-0005,luci2.20181110.0001-0005
  • 5X45 luci1.20181110.0006-0010,luci2.20181110.0006-0010
  • 20X2.5 luci1.20181110.0011-0015,luci2.20181110.0011-0015
  • 25X2.5 luci1.20181110.0016-0020,luci2.20181110.0016-0020
  • 1X2.5 luci1.20181110.0021-0026,luci2.20181110.0021-0026

00:30 TO opens the dome

00:52 AO test

01:20 TO start the collimation on the target Testa A.

01:38 start with the target Testa A
  • acq mods1r.20181110.0001-0004
  • sci mods1r.0005-0007,mods1b.0001-0003
  • seeing 0.95"
  • 02:34 seeing is increasing 1.2"-1.3" ( during the integration of the second image)
  • clouds are coming, the second image is not good
  • 02:40 seeing 1.0'
  • 02:48 seeing very variable (1,0"-1.4") due to clouds that come and go
  • there are too many clouds, we stop the integration, also the last image is not good (we stopped 2 minuts before the end of the script)

03:05 slew to Antoniucci, target V733 Cep
  • acq mods1r.0008-0010
  • sci mods1r.0011-0014, mods1b.0004-0007
  • seeing 1.2" on dimm (the clouds are thinner in this part of the sky )
  • 03:25 seeing about 1.0"-1.1"
  • 03:34 seeing 0.9" 1.0"
  • the clouds are just on the horizont

03:44 slew to standard BD+28
  • acq mods1r.0015-0016
  • sci red grating mods1r.0017-0019
  • seeing 0.9"
  • dual grating mods1r.0020-0022,mods1b.0008-0010
  • 04:04 seeing 0.8

04:14 slew to Fabrizio program,
  • acq mods1r.0023-0025
  • 04:32 execMODS Shall1.ob mods1r.0026,mods1b.0011
  • seeign about 1.00"

05:10 Shall2
  • acq mods1r.0027-0029
  • 05:21 sci mods1r.0030,mods1b.0012
  • seeing about 1.0"
  • 05:32 seeng 1.0-1.1"on the dimm, fwhm= 0.8"-0.9"

05:55 swith LUCI2.

06:09 telluric for Perna HIP16111
  • acq luci2.0026-0028
  • sci --luci2.0029-0032==
  • seeing 1.0"

06:24 slew to target of perna RCSGA327
  • acq =luci2.0033-0038
  • 06:45 send science, but there is an error about detector focus
  • we re-initialize the detector focus, resent the confinguration and sent again the science. the script works luci2.0039-0040
  • SA suggest to stop, and follow again the procedure adding the step to toggle filter back and forth, because after initializing the detector focus, the detector focus does not go to the correct one for the filter+camera combination unless the filter+camera is sent again.
  • 06:58 we follow the procedure
  • 07:00 start again the script. we are not sure the quality of the images luci2.0039-0040
  • luci.0041-0053
  • seeing 0.8"
  • 07:28 SA check the quality of the first images and said that are ok!!!

07:40 slew to Nisini program, field 3, imaging
  • error, loaded the script for both luci
  • send again the script just for luci2. luci2.0054-0079
  • fwhm about 0.6"

08:20 telluric Nisini
  • 08:24 check pointing

08:35 the seeing is good, we slew to AO, target of Pinna HD16760. We take again the data in the filters Ks and H2 to complet the target!!!!The PI said that the first images in H2 and Fe taken two days ago were good.
  • 08:40 send the preset
  • 08:42 error AO
  • 08:48 send again the script
  • 09:00 loop closed, the integration starts luci2.0080-0088 Ks
  • fwhm 0.74, estimated mag 8.9
  • 09:13 dimm 0.9", wfs 0.94", fwhm 0.68", estimated magnitude 9
  • 09:14 SA put the subwindow
  • the SA takes a test image, there is an error in the giers and she can not see the image.
  • 09:28 script sent with the subwindow.
  • 09:29 error.
  • SA try to uderstand and solve the issue operating on LUCI by remote terminal

09:41 TO closes the dome for the humidity, it rices 95%.
  • Temperture at the dew point is about -1.8 C.
  • external temperature is about -0.8 C
  • we will wait until the humidity decreases to less then 90% and remains at that level for almost 30 minutes to open the dome again.

Calibrations

MODS
  • Fabrizio flat + lamp mods.16.Tri2Fab.mms, shall2
  • Fabrizio flat + lamp mods.19.Tri2Fab.mms, shall1
  • Testa flat 1 arcsec mods1r.-0041,mods1b.-0023
  • Testa arc 1 arsec mods1r.0042=0044,mods1b-0024-0026
  • Fabrizio flat + lamp mask ID592936 mods1r.0045-0049, mods1b.0027-0031
  • Fabrizio flat + lamp mask ID551423,deep2 mods1r.0050-0054, mods1b.0032-36
  • Mods2 stuck, we update the gui and everything works
  • Mancini H47-cal-wave-MODS2-blue.txt mods1r.0055-0059, mods1b.0037-0041 no good
  • Mancini H47-MOS-flat-MODS1.txt mods1r.0060-0069, mods1b.0042-0051
  • Mancini slitless mods1r.0070-0079,mods1b.0052-071 no good
Topic revision: r7 - 19 Nov 2018, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback